OpenStudy (anonymous):

Find the inverse function of f(x)= (2-x^3)^5

OpenStudy (raden):

@Coolsector f(x)^(1/5) = 2-x^3 ---> x^3 = 2- f(x)^(1/5) :)

OpenStudy (anonymous):

oops! you are right f(x)^(1/5) = 2-x^3 x^3 =2 - f(x)^(1/5) ill call f-1(x) the inverse of f(x) so : f-1(x) = (2 - x^(1/5) )^(1/3)

OpenStudy (anonymous):

im confused after -f(x)^(1/5)+2=x^3

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

what is my next step

OpenStudy (anonymous):

1. first you want to switch x and y (and since f(x) can also be written as y) for example: y = (2-x^3)^5 (eqn. 1) will become x = (2-y^3)^5 (eqn. 2) Then you want to isolate y from eqn. 2. using algebra

OpenStudy (anonymous):

take the 5th root of both sides

OpenStudy (anonymous):

Yes.. so then you'll have \[\sqrt[5]{x} = (2-y ^{3})\] so what will you do?